# A Simple Latent Diffusion Approach for Panoptic Segmentation and Mask Inpainting

### v1

@@ 3.2 Stage 2: Train a Denoising Diffusion Model
Image-Conditioned Diffusion Process.
Figure 3: Diffusion Process and SNR. (1) During training we randomly sample a timestep from $[1,T]$ in the denoising process. We can increase the RGB image’s importance by strengthening the noise: (i) Following [59, 12], we downscale the latents $z_{c}$ using scaling factor $s\in\mathbb{R}$ and demonstrate its impact – Row 1 ($s=1.0$) is clearly easier to decode than row 2 ($s\approx 0.18$ [59]). (ii) Losses for timesteps near 00 are further downscaled to avoid overfitting. Both strategies enforces the model to focus on the RGB image in generating plausible segmentation maps. Note that we don’t apply explicit constraints to the prior distribution $p(z_{t})$, e.g., match a standard Gaussian $\mathcal{N}(0,1)$. (2) During sampling the denoising process is traversed from right to left in $T_{s}$ iterations.
The second stage of the framework models the function $h$ via a conditional diffusion process. In particular, we aim to learn a diffusion process over discrete timesteps by conditioning the model on images and noisy segmentation masks for each individual timestep. We further follow the formulation of Stable Diffusion [59]. Hence, this process is carried out in a joint latent space – using our trained segmentation encoder $f_{t}$ and given image encoder $f_{i}$ [59] – in order to project the targets and images to their respective latent representations $z_{t}$ and $z_{i}$. We discuss the training and inference procedures next.
@@
where each timestep $j$ is uniformly sampled from $[1,T]$. To reduce training time, we downscale the loss for small timesteps, i.e. $j<25\%\cdot T$. These latents have a high signal-to-noise ratio (SNR) and are thus relatively easy to denoise without modeling semantics. We finally refer to algorithm 1 for an overview of the forward pass during training.
Algorithm 1 Forward pass in Pytorch.
# f_i, f_t: image encoder, segmentation encoder
# h: denoising UNet
@@
y = 2 * y - 1      # bit maps to range [-1, 1]
x = x / 127.5 - 1   # images to range [-1, 1]
z_t = f_t(y) * s    # encode bit maps and rescale
z_i = f_i(x) * s    # encode images and rescale
@@
z_t = scheduler.add_noise(z_t, noise, j)  # Eq. 3
@@
 scheduler.weights: attribute - array of length $T$ with loss weights
Algorithm 1 Forward pass in Pytorch.
Algorithm 2 Sampling process in Pytorch.
# f_i, g: image encoder, segmentation decoder
@@
# T, Ts: number of training steps, inference steps
x = x / 127.5 - 1  # image to range [-1, 1]
z_i = f_i(x) * s    # encode image and rescale
@@
y_pred = g(z_t)  # decode segmentation latents
  scheduler.inference_steps: attribute - chosen sampling timesteps
@@
Algorithm 2 Sampling process in Pytorch.
During sampling, the denoising process is traversed from right to left (see figure 3). It starts from Gaussian noise and progressively adds more details to the segmentation map as controlled by the input image. We rely on the DDIM scheduler [63] to apply this denoising process over a small number of sampling timesteps $T_{s}<<T$. Here, the previous sample is computed as
@@
which follows directly from equation 3. Recall that this strategy allows us to model $p(y|z_{t},z_{i})$, as defined in equation 1. Finally, algorithm 2 provides the details of the sampling process and figure 3 serves as an illustration.
Image Descriptors.
Complementary, we briefly experimented with adding guidance via Stable Diffusion’s cross-attention layers to improve sample quality. This mechanism was initially intended for text embeddings, but can be used to feed any image descriptor. We didn’t observe improvements with self/weakly-supervised priors, like image embeddings from CLIP [55], or when using captions from BLIP [43]. We currently bypass the cross-attention layers with skip connections, but this is a promising direction to explore further.
@@ 3.3 Segmentation Mask Inpainting
Setup.
Our image-conditioned diffusion model is well-suited to complete partial segmentation masks. This is potentially useful for (i) completing sparse 2D segmentation masks obtained from projecting 3D point cloud segmentation labels onto images; (ii) refining noisy segmentation labels that were propagated from one frame to the next using optical flow; (iii) interactive image labeling with rough brush strokes. In contrast to image editing [59], the considered segmentation maps contain empty regions, which we initialize with 00’s. We simulate different sparsities in section 4.2.
Existing state-of-the-art approaches are not designed for these applications as they decouple the output from the input via bipartite matching [16, 8]. They require additional steps to match predictions with the given partial segmentation IDs (e.g., via majority voting). In contrast, diffusion models act on the corrupted masks by default.
Inpainting Process.
Ideally, we can tackle inpainting problems out-of-the-box, i.e., without finetuning. To achieve this goal, the inference loop, previously discussed in algorithm 2, is modified. Assume that we have a dataset of pairs $(y,m)$. Each pair contains a sparse segmentation mask $y\in\{0,1\}^{\log_{2}N\times H\times W}$, represented as a bitmap, and a valid mask $m\in\{0,1\}^{H\times W}$, represented as a boolean mask. Now, the diffusion process should fill in the missing regions in $y$, determined by the zeros in $m$. At each step of the denoising process, the latents corresponding to the given pixels in $m$ are fixed. The key differences with the sampling process are highlighted in algorithm 3.
Algorithm 3 Inpainting process in Pytorch.
# f_i, f_t: image encoder, segmentation encoder
@@
# m: boolean mask indicating valid pixels, size [H, W]
# T, Ts: number of train steps, inference steps
y = 2 * y - 1     # bit maps to range [-1, 1]
@@
x = x / 127.5 - 1  # image to range [-1, 1]
z_i = f_i(x) * s    # encode image and rescale
 z_t_masked = f_t(y) * s  # encode bit map and rescale
@@
y_pred = g(z_t)  # decode segmentation latents
  scheduler.inference_steps: attribute - chosen sampling timesteps
scheduler.remove_noise: function - remove noise from sample
scheduler.step: function - predict previous sample
Algorithm 3 Inpainting process in Pytorch.
@@ 4 Experiments
Figure 4: Qualitative Results on COCO. The figure displays results on COCO val2017. We follow the inference setup (section 3.2) to sample from our model. Only the $\operatorname*{arg\,max}$ operator is applied for post-processing. Key observation: our model disentangles overlapping instances in challenging scenes without complex modules or post-processing. To visualize, segments are assigned to random colors, and missing (VOID) pixels in the ground truth are black.
Figure 5: Qualitative Results on Cityscapes. The figure displays predictions on Cityscapes’s validation set. Key observation: the model produces panoptic masks covering various visual concepts. Thin structures, e.g., poles, are difficult to capture due to the diffusion process in latent space.
@@
Figure 6: Mask Inpainting. The figure visualizes generated samples for different granularity levels by following section 3.3. Key observation: the model is able to fill in missing regions by propagating the partially given segmentation IDs using an image-conditioned diffusion process. Global mask inpainting (left) works out-of-the-box while sparse mask inpainting (right) shows inaccuracies. We hypothesize that this can be addressed by further finetuning LDMseg on sparse inpainting masks.
Figure 7: Ablation Sampling Steps. The figure shows the impact of the number of sampling steps on the PQ, SQ, and RQ metrics (COCO val).
Figure 8: Ablation KL Loss. The figure shows the impact of a KL loss on the PQ, SQ, and RQ metrics (COCO val).
Figure 7: Ablation Sampling Steps. The figure shows the impact of the number of sampling steps on the PQ, SQ, and RQ metrics (COCO val).
Figure 8: Ablation KL Loss. The figure shows the impact of a KL loss on the PQ, SQ, and RQ metrics (COCO val).
Figure 9: Sparsity Experiment. The figure shows the PQ for different drop rates and block sizes ($B\times B$ pixls) in mask $m$ (COCO val).
Figure 10: Impact of Gaussian Noise. We show predictions on COCO val2017 when sampling different noise maps. The images and the generated masks are visualized in row 1 and rows 2 & 3 respectively. All predictions in the same row were generated from the same noise, visualized in column 1. Key observation: different initializations generate panoptic masks with different IDs (see row 2 vs. row 3). The Gaussian noise implicitly functions as object queries [15, 8] or region proposals [29].
4.1 Experimental Setup
Dataset.
We conduct the bulk of our experiments on COCO [45]. The standard train and val sets are adopted for training (118k images) and evaluation (5k images) respectively. During training, we use the panoptic masks – containing both stuff and things classes. For the class-agnostic setup, we only aim to predict instances, without their categories.
Architecture and Training Setup.
@@ 4.1 Experimental Setup
The maximum number of detectable segments $N$ is set to $128$. We resize the input to $512\times 512$, apply random horizontal flipping, and randomly assign segmentation IDs to integers in [$10,N-$ 1]. The segmentation encoder $f_{t}$ processes the panoptic mask $y$ by leveraging 3 convolutional layers with stride 2 and SiLU [22] activations. This results in a resizing factor $f$ of $8$ and latents with size $4\times 64\times 64$. We rely on the image encoder $f_{i}$ from Rombach et al. [59] (VAE) to convert the image $x$ to latents of the same size. We adopt Stable Diffusion’s pretrained weights to initialize the UNet $h$, and its rescaling factor $s$ to lower the SNR as $s\cdot z_{c}$ [59]. Notably, $4$ zero-initialized channels are appended to the first convolutional layer of $h$, allowing us to operate on the concatenated input $z_{c}$. We additionally incorporate self-conditioning [13] to improve sample quality. Our segmentation decoder $g$ consists of 2 transpose convolutions, resulting in an upscaling factor of $4$. We randomly sample $j$ from $1000$ discrete timesteps and linearly decay the loss for the bottom $25$% to lower the impact of samples with high SNR. The AdamW [48] optimizer is adopted with a learning rate of $1e^{-4}$ and weight decay of $5e^{-2}$. Finally, the first stage is trained for 60k iterations with a batch size of $16$ while the second stage is trained for 50k iterations with a batch size of 256 unless stated otherwise.
Inference Setup and Evaluation Protocol.
@@
During inference, the DDIM scheduler [64] generates samples with $50$ equidistant timesteps in latent space. At the end of the denoising process, we decode and upscale the segmentation logits (output of $g$) with a factor of $2$ using bilinear interpolation. The $\operatorname*{arg\,max}$ operator produces the final (discrete) per-pixel segmentation masks. We benchmark our approach with the Panoptic Quality (PQ) evaluation protocol, defined by Kirillov et al. [38]. PQ is the product of two quality metrics: the segmentation quality which measures the intersection-over-union of matched segments (IoU) and the recognition quality which measures the precision and recall (F-score).
4.2 Panoptic Segmentation Results
Image-conditioned Mask Generation.
We visualize predictions of LDMSeg in figure 4 and 5 on a few validation images from COCO [45] and Citysapes [19] respectively. The model divides an image into semantically meaningful groups and captures different objects in the scene, e.g., persons, horses, cats, cars, etc. Notably, LDMSeg directly generates non-overlapping instance masks, required for panoptic segmentation. Additionally, figure 10 visualizes the predictions when sampling from our generative model (see first column) with different seeds. Different Gaussian noise maps lead to different segmentation IDs. Compared to other frameworks, the sampled noise resembles the object queries in Mask2Former and DETR [16, 8], or the regions proposals in Mask R-CNN [29].
Mask Inpainting.
Figure 6 demonstrates the inpainting performance for two different granularity levels in mask $m$. We mask random regions in the ground truth to simulate global and local completion tasks. The results demonstrate that LDMSeg is able to fill panoptic masks via the diffusion process, and without requiring additional components.
@@ 4.2 Panoptic Segmentation Results
Further, figure 9 displays the PQ metric for a wide range of drop rates $d$ and block sizes $B$. While our model performs well for global inpainting, there is still room for improvement when considering very sparse inpainting tasks (e.g., dropping $16\times 16$ pixels with a drop rate of $90\%$). Particularly, for drop rates above $70\%$ and small block sizes, the reduction rate (negative slope) in PQ increases. Further improvements can be obtained by finetuning or modifying the autoencoder. For instance, shallower encoders $f_{t}$ could adapt better to sparse inputs, in exchange for lower reconstruction quality.
  Method Backbone PQ $\uparrow$
@@
  ${}^{\dagger}$ (Ours) UNet [60] 50.9
   LDMSeg (Ours) UNet [60] 50.8
Table 1: Comparison with Specialists (class-agnostic). The results are reported on COCO val2017. We differentiate between specialist and generalist methods and train for 50k iterations. Training details for Mask2Former are provided in the supplementary since we use a simplified version, similar to SAM [40] with MAE [30] or DINOv2 [53] initialization (see normal-$\star$). (normal-$\dagger$) denotes the adoption of Stable Diffusion’s VAE with 84M trainable parameters, resulting in 20×20$\times$20 × more parameters compared to our standard setup (84M vs. 2M) for the same performance (see section 3.1).
  Method Backbone PQ $\uparrow$
@@
  MaskFormer [15] ResNet [28] 46.5
  Mask2Former [16] Swin-L [46] 57.8
@@
  Painter [78] ViT [21] 41.3
  UViM [41] ViT [21] 45.8
  Pix2Seq-$\mathcal{D}^{\dagger}$ [12] ResNet [28] 50.3
   LDMSeg (Ours) UNet [60] 43.3
Table 2: State-of-the-art Comparison. The table presents the panoptic segmentation results on COCO val2017. (LDMSeg vs. specialists) Our approach almost matches the performance of several specialists on the panoptic segmentation task. (LDMSeg vs. generalists) LDMSeg outperforms Painter, but not Pix2Seq-$\mathcal{D}$ and UViM. Note that Pix2Seq-$\mathcal{D}$ leverages additional annotated training data (normal-$\dagger$) from Objects365 [61] and UViM is sensitive to the adopted guiding codes, necessitating special mechanisms like code dropout and optimal code length [41]. In contrast, we can perform mask inpainting out-of-the-box.
State-of-the-art Comparison.
Table 1 compares LDMSeg with prior art for the class-agnostic setup. Our model obtains 50.8% PQ on COCO’s validation set. However, there is still a gap with specialized methods, i.e., MaskFormer [15, 16] which obtains 59.0% PQ.
To compare with state-of-the-art methods, we train our method for 200k iterations, as well as a classifier with a simple segmentation head [9] on top of the proposals. We compare the obtained panoptic segmentation in table 2. Our generative approach can almost match several specialized approaches ($43.3\%$ vs. $44.1\%$[79, 37]. LDMSeg outperforms Painter [78] while not requiring different encoding strategies for things and stuff categories or NMS ($43.3\%$ vs. $41.3\%$). While UViM [41] obtains better results, it’s arguably less robust due to its strong reliance on specific code lengths and code dropout, in order to train its autoregressive language model (LM). In contrast, we leverage latent codes centered around zero using a shallow autoencoder, which benefits robustness. To increase the denoising difficulty, we simply drop the scaling factor $s$ (see section 3.2). Unsurprisingly, Pix2Seq-$\mathcal{D}$ [12] achieves higher PQ by adopting additional training data [61] and a pixel-level objective. We expect that larger datasets and a higher resolution will help in closing this gap.
In contrast to prior art [41, 78], LDMSeg can tackle the task of mask inpainting out-of-the-box, and without specific processing techniques. As UViM [41] is not capable of mask inpainting by design, it’s unclear how Painter [78] and Pix2Seq-$\mathcal{D}$ [12] should be modified: (i) Painter [78] relies on a predefined color palette that differs for things and stuff classes. Due to this fixed encoding (e.g., center of mass for things categories), it is unable to complete the given (random) segmentation IDs. Furthermore, its hefty post-processing is not guaranteed to preserve the IDs in the partial input mask, which is detrimental for mask inpainting. Our method is conditioned on latents that are not confined to a task-specific encoding, which preserves the given input masks. (ii) Future research could explore if our inpainting strategy (see section 3.3) can be extended to Bit Diffusion [13], used in Pix2Seq-$\mathcal{D}$ [12]. However, the unreleased weights poses a challenge in adopting this paradigm.
Ablation Studies.

### v2

@@ 3.2 Stage 2: Train a Denoising Diffusion Model
Figure 3: Diffusion Process and SNR. (1) During training we randomly sample a timestep from $[1,T]$ in the denoising process. We can increase the RGB-image’s importance by strengthening the noise: (i) Following [64, 10], we downscale the latents $z_{c}$ using scaling factor $s\in\mathbb{R}$ and demonstrate its impact – Row 1 ($s=1.0$) is clearly easier to decode than row 2 ($s\approx 0.18$ [64]). (ii) Losses for timesteps near 00 are further downscaled to avoid overfitting. Both strategies force the model to focus on the RGB image in generating plausible segmentation maps. Note, we don’t apply explicit constraints to the prior distribution $p(z_{t})$, e.g., match a standard Gaussian $\mathcal{N}(0,1)$. (2) During sampling the denoising process is traversed from right to left in $T_{s}$ iterations.
Image-Conditioned Diffusion Process. The second stage of the framework models the function $h$ via a conditional diffusion process. We aim to learn a diffusion process over discrete timesteps by conditioning the model on images and noisy segmentation masks for each individual timestep. We further follow the formulation of Stable Diffusion [64]. This process is carried out in a joint latent space – using our trained segmentation encoder $f_{t}$ and given image encoder $f_{i}$ [64] – in order to project the targets and images to their respective latent representations $z_{t}$ and $z_{i}$. Next, we will discuss the training and inference procedures.
@@
where each timestep $j$ is uniformly sampled from $[1,T]$. To reduce training time, we downscale the loss for small timesteps, i.e. $j<25\%\cdot T$. These latents have a high signal-to-noise ratio (SNR) and are thus relatively easy to denoise without modeling semantics. We refer to Algorithm 1 for an overview of the forward pass.
# f_i, f_t, h: encoders f_i & f_t, UNet h
@@
y = 2 * y - 1,
x = x / 127.5 - 1
z_t = f_t(y) * s
z_i = f_i(x) * s
@@
# apply Eq. 2
z_t = scheduler.add_noise(z_t, noise, j)
@@
Algorithm 1 Forward pass.
 scheduler.weights: array of length $T$ with loss weights
# f_i, g: image and segmentation decoders
@@
# T, Ts: #training steps, #inference steps
x = x / 127.5 - 1
z_i = f_i(x) * s
@@
y_pred = g(z_t) # decode latents
Algorithm 2 Sampling process.
  scheduler.inference_steps: chosen sampling timesteps
@@
During sampling, the denoising process is traversed from right to left (see Figure 3). It starts from Gaussian noise and progressively adds more details to the segmentation map as controlled by the input image. We can rely on the DDIM scheduler [68] to apply this denoising process over a small number of sampling timesteps $T_{s}<<T$. Here, the previous sample is computed as
@@
which follows directly from Equation 2. Recall that this strategy allows us to model $p(y|z_{t},z_{i})$. Finally, Algorithm 2 provides the details of the sampling process and Figure 3 serves as an illustration.
Image Descriptors. Complementary, we briefly experimented with adding guidance via Stable Diffusion’s cross-attention layers to improve sample quality. This mechanism was initially intended for text embeddings, but can be used to feed any image descriptor. We didn’t immediately observe improvements with self/weakly-supervised priors, like image embeddings from CLIP [60], or when using captions from BLIP [45]. Hence, we can bypass the cross-attention layers with a skip connection in the class-agnostic setup.
Task Embeddings. To extend our framework to multiple tasks, we add learnable embeddings. We query the model for a certain task via its cross-attention layers. In particular, (class-aware) panoptic masks can be obtained by merging the semantic and instance predictions, which necessitates two task embeddings.
@@ 3.3 Segmentation Mask Inpainting
Setup. Our image-conditioned diffusion model is well-suited to complete partial segmentation masks. This is potentially useful for completing sparse segmentation masks obtained from projected point clouds or for interactive image labeling with rough brush strokes. In contrast to image editing [64], the considered segmentation maps contain empty regions, which we initialize with 00’s. We simulate different sparsities in Section 6. Existing state-of-the-art approaches are not designed for these applications as they decouple the output from the input via bipartite matching [15, 9]. They require additional steps to match predictions with the given partial segmentation IDs (e.g., via majority voting). In contrast, diffusion models act on the corrupted masks by default.
# f_i, f_t: image encoder, segm. encoder
@@
# m: boolean mask w/ valid pixels, size [H, W]
# T, Ts: #train steps, #inference steps
y = 2 * y - 1
@@
x = x / 127.5 - 1
z_i = f_i(x) * s
 z_t_masked = f_t(y) * s
@@
y_pred = g(z_t) # decode latents
Algorithm 3 Inpainting process.
Inpainting Process. Ideally, we can tackle inpainting problems out-of-the-box, i.e., without finetuning. To achieve this goal, the inference loop, previously discussed in Algorithm 2, is modified. Assume that we have a dataset of pairs $(y,m)$. Each pair contains a sparse segmentation mask $y\in\{0,1\}^{\log_{2}N\times H\times W}$, represented as a bit map, and a valid mask $m\in\{0,1\}^{H\times W}$, represented as a boolean mask. Now, the diffusion process should fill in the missing regions in $y$, determined by the zeros in $m$. At each step of the denoising process, the latents corresponding to the given pixels in $m$ are fixed. The key differences with the sampling process are highlighted in Algorithm 3.
@@ 4 Experiments
Figure 4: Qualitative Results. The figure displays results on COCO val2017. We follow the inference setup (Section 3) to sample from our model. Only the $\operatorname*{arg\,max}$ operator is applied for post-processing. Our model disentangles overlapping instances in challenging scenes without complex modules or post-processing. To visualize, segments are assigned to random colors, and missing (VOID) pixels in the ground truth are black.
@@
Figure 5: Mask Inpainting. The figure visualizes generated samples for different granularity levels by following Section 3.3. The model can fill in missing regions by propagating the partially given (random) segmentation IDs using an image-conditioned diffusion process. Global mask inpainting (left) results are reasonable out-of-the-box while sparse mask inpainting (right) shows inaccuracies. We hypothesize that this can be addressed by further finetuning LDMseg on sparse inpainting masks.
Dataset. We conduct the bulk of our experiments on COCO [47] by relying on the panoptic masks with stuff and things classes. It contains 118k and 5k images for training and evaluation respectively.
@@
Architecture and Training Setup. The maximum number of detectable segments $N$ is set to $256$. We resize the input to $512\times 512$, apply random horizontal flipping, and randomly assign integers from [$10,N-$ 1] to the segments. The segmentation encoder $f_{t}$ processes the panoptic mask $y$ by leveraging 3 convolutional layers with stride 2 and SiLU [22] activations. This results in a resizing factor $f$ of $8$ and latents with size $4\times 64\times 64$. We rely on the image encoder $f_{i}$ from Rombach et al. [64] (VAE) to convert the image $x$ into latents. We adopt Stable Diffusion’s pretrained weights to initialize the UNet $h$, and its rescaling factor $s$ to lower the SNR as $s\cdot z_{c}$ [64]. Notably, $4$ zero-initialized channels are appended to the first convolutional layer of $h$, allowing us to operate on the concatenated input $z_{c}$. Further, self-conditioning [13] is used to improve sample quality. Our segmentation decoder $g$ consists of 2 transpose convolutions, resulting in an upscaling factor of $4$. We randomly sample $j$ from $1000$ discrete timesteps and linearly decay the loss for the bottom $25$% to lower the impact of samples with high SNR. The AdamW [50] optimizer is adopted with a learning rate of $1e^{-4}$ and weight decay of $1e^{-1}$. Finally, the first stage is trained for 60k iterations with a batch size of $8$ while the second stage is trained for 50k iterations with a batch size of 256 unless stated otherwise.
@@
Inference Setup and Evaluation Protocol. During inference, the DDIM scheduler [70] generates samples with $50$ equidistant timesteps in latent space. At the end of the denoising process, we decode and upscale the segmentation logits (output of $g$) with a factor of $2$ using bilinear interpolation. The $\operatorname*{arg\,max}$ operator produces the final (discrete) per-pixel segmentation masks. We benchmark our approach with the Panoptic Quality (PQ) evaluation protocol, defined by Kirillov et al. [40]. PQ is the product of two quality metrics: the segmentation quality which measures the intersection-over-union of matched segments (IoU) and the recognition quality which measures the precision and recall (F-score).
4.1 Segmentation Results
Image-conditioned Mask Generation. The model directly divides an image into non-overlapping semantic masks, capturing different objects in the scene, such as persons, horses and cars. Figure 4 showcases the predictions.
Figure 6: Impact of Gaussian Noise.
As our model is inherently stochastic, we show the predictions when sampling different seeds in Figure 6 (first column). Rows 2 and 3 show the predictions when starting from the same noise map. The key observation is that different maps generate different segmentation IDs. Compared to other frameworks, the sampled noise resembles the object queries in Mask2Former or DETR [15, 9], and the regions proposals in Mask R-CNN [29]. Table 1 quantitatively compares LDMSeg with prior art for the class-agnostic setup. We obtain 50.8% PQ on COCO val. However, there is still a gap with specialized methods, e.g., Mask2Former [15] reaches 59.0% PQ.
@@ 4.1 Segmentation Results
Mask Inpainting. Next, we mask random regions in the ground truth to simulate global and local completion tasks. Figure 5 shows the inpainting performance for two different granularity levels in mask $m$. The results demonstrate that LDMSeg is able to complete panoptic masks via the diffusion process, and without requiring additional components. Figure 9 displays the PQ metric for a wide range of drop rates $d$ and block sizes $B$. While our model is able to complete global regions, there is still room for improvement when considering sparse inpainting tasks (e.g., dropping $16\times 16$ pixels with a drop rate of $90\%$). For drop rates above $70\%$ and small block sizes, the reduction rate (negative slope) in PQ increases. Further improvements can be obtained by finetuning or modifying the autoencoder. For instance, shallower encoders $f_{t}$ could adapt better to sparse inputs, in exchange for lower reconstruction quality. Others [79, 43, 15] are not designed to complete partial masks without complex post-processing.
Table 1: Class-agnostic Comparison. We compare with Mask2Former on COCO val2017. Training details for Mask2Former are specified in Supplement A, following SAM [41] with MAE [28] or DINOv2 [58] ($\star$) initialization. ($\dagger$) denotes the adoption of Stable Diffusion’s VAE with 84M parameters, resulting in 40×40$\times$40 × more parameters (84M vs. 2M) for the same performance (Section 3.1). $PQ_{inpaint}$ is the averaged PQ metric when dropping $16\times 16$ pixels with probabilities from 10 to 90%.
     Method Backbone PQ $\uparrow$ PQinpaint $\uparrow$
@@
     LDMSeg UNet [65] 50.9
   LDMSeg UNet [65] 50.8 61.3
Table 2: State-of-the-art Comparison. The table presents the panoptic and semantic segmentation results on COCO val2017 and ADE20k val respectively.
COCO [47] ADE20k [91]
     Method Backbone PQ $\uparrow$ mIoU $\uparrow$
@@
     MaskFormer [16] ResNet [30] 46.5 44.5
     UPerNet [84] Swin-L [48] 52.1
     DDP [34] Swin-L [48] 53.2
     Mask2Former [15] Swin-L [48] 57.8 56.1
@@
     Painter [79] ViT [21] 41.3 47.3
     UViM [43] ViT [21]           43.1 (45.8)
   LDMSeg UNet [65] 44.3 52.2
Table 3: Multi-Task Performance. The table reports results for semantic segmentation, panoptic segmentation and relative depth (between 0 - 1) on COCO val2017.
Semantic Seg. Panoptic Seg. Relative Depth
Setting # Iters mIoU $\uparrow$ PQ $\uparrow$ RMSE $\downarrow$
Multi-Task 50 60.1 44.1 0.075
Single-Task 50 61.5 44.3 0.067
State-of-the-art Comparison. As the panoptic task combines instance and semantic segmentation, we introduce two learnable task embeddings. Additionally, we adopt a ViT-B [21] as the image encoder to improve the performance (see Supplement B for a component analysis). Table 2 compares LDMSeg with state-of-the-art methods after training for 100k iterations:
  • LDMSeg vs. Specialists: Our generative approach is competitive with several specialized approaches [85, 39, 9]. For instance, we can match the performance of PanopticFPN [40] while not requiring region proposals. Instead, we rely on the diffusion process to solve the permutation-invariance of the instances.
  • LDMSeg vs. Generalists: (i) Painter necessitates two separate encoding schemes for things and stuff categories, as well as non-maximum-suppression (NMS). LDMSeg outperforms Painter ($44.3\%$ vs. $41.3\%$) while not relying on these components, nor on complex merging strategies. This renders our approach more general. In fact, our post-processing time is an order of magnitude smaller (see Supplement B). (ii) Additionally, LDMSeg is competitive with UViM’s [43] public model ($44.3\%$ vs. $43.1\%$). We hypothesize that the disparities in UViM’s results could be attributed to its reliance on specific code lengths and code dropout [43], in order to train its autoregressive language model (LM). In contrast, we leverage latent codes centered around zero using a shallow autoencoder, enabling better control. To increase the denoising difficulty, we simply lower the scaling factor (see Section 3.2). Pix2Seq-$\mathcal{D}$ [10] relies on a ResNet backbone that is pretrained with additional bounding-box annotations (Objects365 [66]) and achieves $50.3\%$ PQ. However, its unreleased weights poses a challenge in adopting this paradigm. To our knowledge, LDMSeg is first in demonstrating that a latent diffusion process can bypass object detection data and its related modules. Larger datasets and a higher resolution will help in closing the gap.
Finally, we tackle semantic segmentation and show results for ADE20k [91]. This dataset contains 20k training images and 2k validation images, covering 150 semantic classes. LDMSeg surpasses well-performing methods [79, 16, 84], reaching 52.2% mIoU on the validation set. Interestingly, our model is able to capture the mapping from latents to the respective classes without making changes to its design. While DDP [34] performs slightly better (52.2% vs. 53.2%), it can not handle the permutation-invariance of the instance masks. This limitation stems from relying on the diffusion process as a refinement step and its necessity for a direct mapping between inputs and outputs (e.g., in semantic segmentation).
4.2 Multi-Task Learning Results
We now broaden the scope and extend our approach to a multi-task setting. Table 3 reports the results when training on 3 vision tasks for 100k iterations on COCO in total. We handle all tasks identically, including the same scaling factor and augmentations. Given that the single-task setup has been trained for 3 $\times$ longer, it serves as an upperbound (e.g., 44.2% vs. 44.3% for PQ). Most importantly, we conclude that LDMSeg can be trained on multiple dense prediction tasks simultaneously by leveraging task embeddings. We refer to Supplement A for more details and visualizations.
Figure 7: Sampling Steps. The impact of sampling steps on PQ, SQ, and RQ.
Figure 8: KL Loss. The impact of a KL loss on PQ, SQ, and RQ.
Figure 9: Mask Inpainting. PQ for varying drop rates and block sizes.
4.3 Ablation Studies
We ablate LDMSeg in the class-agnostic setting after training for 50k iterations.